When drawing cards without replacement from a standard 52- card deck, find the maximum number of cards you could possibly draw and still get
(a) fewer than three black cards,
(b) fewer than six spades,
(c) fewer than four face cards,
(d) fewer than two kings

Answer by richwmiller(17219)

a) all the red cards +2 black
26+2=28
b)all the non spades
39+5=44
c)there are 4*3=12 face cards
40 non face cards
40+3=43
d)48+1=49